[Hallc_running] RC update 04/04/2019

gabriel at jlab.org gabriel at jlab.org
Thu Apr 4 08:43:51 EDT 2019

Dear all-
Here is a quick RC update:
- no major problems to report. A SHMS Q1 reset at the beginning of swing
shift (6-10 min lost) and a transitory alarm (no downtime) on the raster
during the owl shift.
- Smooth/uneventful running for the swing shift 04/03 and owl shift 04/04.
Finished the Q2 = 3.0 GeV^2 low epsilon point at 6.91 degrees with ~7100
good Lambdas (slightly lower than the "wanted" 8200 but there might be
some undercounting - see hclog note about user_ana not going to the
physical end of a run in progress).
- Moved to the "side" setting, \theta_{HMS} = 9.91 degrees this morning.
We will be at this kinematic until Sunday afternoon/evening.
- a BCM calibration is scheduled for this morning (04/04/2019).

Happy Kaon hunting!
Gabriel Niculescu

More information about the Hallc_running mailing list